OpenStudy (moonlitfate):

Write the expression in algebraic form. [Hint: Sketch a right triangle.] csc (arctan (x/√2))

myininaya (myininaya):

To do questions like these, start with the inside funtion Let u=arctan(x/sqrt(2)) so tan(u)=x/sqrt(2) Draw a right triangle and use this equation Then find the remaining side then when use that triangle to find csc(u) you are done. Please start with the drawing. Then use Pythagorean thm to find missing side. Let me know if you have trouble.

OpenStudy (moonlitfate):

I went and set: \[y= \arctan(\frac{ x }{ \sqrt{2}}\]

OpenStudy (moonlitfate):

And take the inverse of both sides, so tan(y) = \[\frac{ x }{ \sqrt{2} }\]

myininaya (myininaya):

Ok draw a right triangle.

myininaya (myininaya):

Then use that equation to label the side measurements and angle of the triangle

OpenStudy (anonymous):

you get the best people to help you here !

OpenStudy (moonlitfate):

 I hope, I did that right?

myininaya (myininaya):

tan is the ratio that gives you opp/adj

myininaya (myininaya):

\[\frac{x}{\sqrt{2}}=\frac{opp}{adj}\]

OpenStudy (moonlitfate):

Yeah, I know that much. :)

myininaya (myininaya):

I would simply let the opp measurement be x and the adjacent measurement be sqrt(2)

OpenStudy (moonlitfate):

Oh, okay.

myininaya (myininaya):

Do you see how I got that from the equation you gave?

OpenStudy (moonlitfate):